# Revisiting 3D LLM Benchmarks: Are We Really Testing 3D Capabilities?

### v2

@@ Abstract
In this work, we identify the “2D-Cheating” problem in 3D LLM evaluation, where these tasks might be easily solved by VLMs with rendered images of point clouds, exposing ineffective evaluation of 3D LLMs’ unique 3D capabilities. We test VLM performance across multiple 3D LLM benchmarks and, using this as a reference, propose principles for better assessing genuine 3D understanding. We also advocate explicitly separating 3D abilities from 1D or 2D aspects when evaluating 3D LLMs. Code and data are available at https://github.com/LLM-class-group/Revisiting-3D-LLM-Benchmarks.
@@ 1 Introduction
Recent advances in Large Language Models OpenAI (2022, 2023) have led to the development of Vision-Language Models (VLMs) Liu et al. (2023); OpenAI (2024). To overcome their lack of grounding in the real 3D physical world Hong et al. (2023), researchers developed 3D LLMs Qi et al. (2024); Xu et al. (2023) for 3D processing.
@@
Specifically, we render point clouds into images to test the performance of VLMs on 3D LLM benchmarks. Experimental results indicate that VLMs can significantly exceed the state-of-the-art performance of 3D models on certain benchmarks. However, on other benchmarks, even with enhanced information, they still fail to achieve the best performance.
@@ 2.2 Viewpoint Selection
Object point clouds are relatively simple, so we use a single fixed viewpoint for each benchmark, typically the front or side view of most objects.
@@
Scene point clouds present greater complexity. To systematically investigate viewpoint impacts, we conducted the following experimental settings, as all viewpoints illustrated in Figure 3.
@@ 2.2.1 Single View
In this setting, we follow the common practice Ma et al. (2022) to render images from BEV.
@@ 2.2.3 Oracle View
In this setting, we conducted experiments on the validation set to explore the upper limits of VLM3D’s capabilities. We first rendered images from all five viewpoints shown in Figure 3, and then used the Best-of-N (BoN) method to select the optimal view for each question.
@@
Table 1: Results on object benchmarks. The 3D SOTA models are ShapeLLM-13B and GPT4Point, respectively.
Since LLMs are probabilistic, simply taking the highest evaluation score risks inflating the likelihood of randomly guessing the correct answer. To address this, we sampled multiple responses for each viewpoint and calculated the average score. The highest average score among all viewpoints was then selected as the oracle view score.
2.3 Other Rendering Factors
While other factors like rendering style may affect image quality, viewpoint selection remains the trickiest due to the difficulty of a unified optimal setup. For each benchmark, we applied a consistent set of other rendering configurations across all point clouds to ensure high-quality results.
3 Experiments
We conducted experiments on the benchmarks in Figure 4, which covered major benchmarks used for evaluating 3D LLMs.
3.1 Object Point Cloud Benchmark
For object point cloud benchmarks, we use 3D MM-Vet Qi et al. (2024) for Q&A and ObjaverseXL-LVIS Caption Deitke et al. (2022) for captioning, with results presented in Table 1.
@@ 3.2 Scene Point Cloud Benchmark
We conducted experiments on two widely used scene point cloud benchmarks, ScanQA Azuma et al. (2021) and SQA3D Ma et al. (2022). ScanQA evaluates general Q&A, while SQA3D further assesses situation understanding. We chose the 3D Baseline as the best-performing model at the time of the benchmark’s release.
@@ 3.2.1 Single View Evaluation
We first conducted the single view evaluation mentioned in Section 2.2.1, with results in Table 2.
@@
Table 2: Single-view results: For ScanQA, baseline is ScanQA, SOTA is LEO Huang et al. (2023). For SQA3D, baseline is SQA3D, SOTA is SIG3D Man et al. (2024). Red and green indicate decrease and increase relative to the 3D baseline, respectively.
We observed that these benchmarks often focus on specific objects within intricate scenes, with single view images frequently missing required information. To further explore the upper bound of VLMs, we conducted experiments using different viewpoints on Qwen2-VL-72B Wang et al. (2024).
@@ 3.2.2 Multi View Evaluation
We conducted multi-view evaluation based on the settings in Section 2.2.2, with results in Table 3.
@@
Table 3: The result of multi-view. Baseline and SOTA models are the same in Table 2. SV refers to single-view, and MV refers to multi-view.
@@ 3.2.3 Oracle View Evaluation
We performed oracle view evaluation using the Best-of-N method in Section 2.2.3. For each question, we rendered images from $k$ viewpoints and sampled $n$ responses per viewpoint. As shown in Figure 5, the result decreases as $n$ increases and converges around $n=20$, indicating that randomness from selecting the maximum score is effectively mitigated. The scores are detailed in Table 4.
Besides the Best-of-N, we explored another method called Human-Intuition-Selection (HIS). We calculate the centroids of all relevant objects of each question and use a heuristic algorithm to select the best viewpoint that captures the centroids effectively. Details and results are in Appendix B.1.
@@ 4 Analysis
Results reveal that VLM outperforms 3D LLMs using only single view images on these object benchmarks. This suggests that these tasks can be easily solved without specialized 3D representations, indicating a lack of effective 3D capability assessment. The limitation may arise from (1) $\bullet$ the inherent simplicity of these object point clouds, and (2) $\bullet$ these tasks primarily require only surface-level information that could be equally extracted from a simple rendered image. See more task examples in Appendix A.1 and Appendix A.2.
@@ 4.1.1 Poor Performance with Single View
In these scene benchmarks, even the most advanced VLMs consistently underperform top 3D models, $\bullet$ revealing their limitations in handling complex 3D scenes.
@@ 4.1.4 Failure of HIS Method
Table 5 shows that HIS performs significantly worse than Best-of-N. Through case studies, we identify three key factors: (1) $\bullet$ Selecting the best viewpoint for detail items in complex scenes with intuition is challenging, often leading to occlusions. (2) $\bullet$ Limited ground truth reveals a clear preference, as shown in Appendix C.2. While multiple viewpoints all provide sufficient information, BoN selects the one that aligns most closely with the ground truth. (3) $\bullet$ Many questions rely on common sense, not the specific details of the point cloud, as shown in Appendix C.3. Specific information in HIS viewpoints might even lead to a worse score compared to answering leveraging world knowledge with an unrelated viewpoint in BoN.
@@ 5 Principles For Effective 3D Evaluation
Based on the analysis, we propose the following principles for effective 3D capabilities evaluation of 3D LLMs. [^1] The analyses leading to each principle are marked with corresponding colored circular markers in the preceding text. We applied these principles to redesign a task, as shown in Figure 6.
@@ Acknowledgement
We would like to express our sincere gratitude to Prof. Yonglu Li for his valuable guidance and support throughout this research, from topic selection to the final writing. His insightful discussions and feedback have been essential to the completion of this work. We would also like to thank Ye Wang for kindly sharing the viewpoint dataset in ScanNet.
@@ C.2.2 ScanQA
Here is an example where the Best-of-N method achieved a significantly better score due to its alignment with ground truth preference: the question is “Where is the brown chair located?” and the ground truths are “under the circular table” and “next to the left table”. The answer provided by the Best-of-N method is “on the right side of the table”, based on the input image shown in Figure 14 (a). The Human-Intuition-Selection method gives the answer “on the right side of the table”, using the input image shown in Figure 14 (b).
It is evident that there are many reasonable answers to this question; however, the ground truth prefers responses related to the table. While both viewpoints depict the chair, the answer from the Human-Intuition-Selection method is influenced by the nearby cabinet in its chosen viewpoint, resulting in a lower score.
(a) The viewpoint selected by Best-of-N.
(b) The viewpoint selected by Human-Intuition-Selection.

### v3

@@ Abstract
In this work, we identify the “2D-Cheating” problem in 3D LLM evaluation, where these tasks might be easily solved by VLMs with rendered images of point clouds, exposing ineffective evaluation of 3D LLMs’ unique 3D capabilities. We test VLM performance across multiple 3D LLM benchmarks and, using this as a reference, propose principles for better assessing genuine 3D understanding. We also advocate explicitly separating 3D abilities from 1D or 2D aspects when evaluating 3D LLMs. Code and data are available at Github.
@@ 1 Introduction
Recent advances in Large Language Models OpenAI (2022, 2023) have led to the development of Vision Language Models (VLMs) Liu et al. (2023); OpenAI (2024). To overcome their lack of grounding in the real 3D physical world Hong et al. (2023), researchers developed 3D LLMs Qi et al. (2024); Xu et al. (2023) for 3D processing.
@@
Specifically, we render point clouds into images to test the performance of VLMs on 3D LLM benchmarks. Experimental results indicate that VLMs can significantly exceed the state-of-the-art (SOTA) performance of 3D models on certain benchmarks. However, on other benchmarks, even with enhanced information, they still fail to achieve the best performance.
@@ 2.2 Viewpoint Selection
Object point clouds are relatively simple, so we use a fixed viewpoint for each benchmark. Following common practice, the camera is positioned at a fixed location looking toward the origin of the point cloud. No specific adjustments are made for individual point clouds.
@@
Scene point clouds present greater complexity. To systematically investigate viewpoint impacts, we conduct the following experimental settings, as all viewpoints illustrated in Figure 3.
@@ 2.2.1 Single View
In this setting, we follow the common practice Ma et al. (2022) to render images from Bird’s Eye View (BEV).
@@ 2.2.3 Oracle View
In this setting, we conduct experiments on the validation set to explore the upper limits of VLM3D’s capabilities. We first render images from all five viewpoints shown in Figure 3, and then use the Best-of-N (BoN) method to select the optimal view for each question.
@@
Since LLMs are probabilistic, simply taking the highest evaluation score risks inflating the likelihood of randomly guessing the correct answer. To address this, we sample multiple responses for each viewpoint and calculate the average score. The highest average score among all viewpoints is then selected as the oracle view score.
2.3 Other Rendering Factors
While other factors like rendering style may affect image quality, viewpoint selection remains the trickiest due to the difficulty of a unified optimal setup. For each benchmark, we apply a consistent set of other rendering configurations across all point clouds to ensure high-quality results.
3 Experiments
3.1 Setup
We conduct experiments on the benchmarks in Figure 4, covering major benchmarks used for evaluating 3D LLMs. We compare VLM3D’s performance with current SOTA 3D LLMs: ShapeLLM-13B on 3D MM-Vet Qi et al. (2024), GPT4Point on ObjaverseXL-LVIS Caption Deitke et al. (2022), LEO Huang et al. (2023) on ScanQA Ma et al. (2022), and SIG3D Man et al. (2024) on SQA3D Ma et al. (2022). For scene point cloud benchmarks, we also compare against the best-performing model at the time of the benchmark’s release as 3D Baseline. For the VLM used in VLM3D, we choose GPT-4o OpenAI (2024) and Qwen2-VL-72B Wang et al. (2024).
3.2 Object Point Cloud Benchmark
For object point cloud benchmarks, we choose 3D MM-Vet for Q&A and ObjaverseXL-LVIS Caption for captioning, with results presented in Table 1.
@@ 3.2 Object Point Cloud Benchmark
Table 1: Results on object benchmarks.
@@ 3.3 Scene Point Cloud Benchmark
We conduct experiments on two widely used scene point cloud benchmarks, ScanQA and SQA3D. ScanQA evaluates general Q&A, while SQA3D further assesses situation understanding.
@@ 3.3.1 Single View Evaluation
We first conduct the single view evaluation mentioned in Section 2.2.1, with results in Table 2.
@@
Table 2: Result for single-view evaluation on scene benchmarks. Red and green indicate decrease and increase relative to the 3D Baseline, respectively.
We observe that these benchmarks often focus on specific objects within intricate scenes, with single view images frequently missing required information. To further explore the upper bound of VLMs, we conduct experiments using different viewpoints on Qwen2-VL-72B.
@@ 3.3.2 Multi View Evaluation
We conduct multi-view evaluation based on the settings in Section 2.2.2, with results in Table 3.
@@
Table 3: The result of multi-view evaluation on scene benchmarks. SV refers to single-view, and MV refers to multi-view.
@@ 3.3.3 Oracle View Evaluation
We perform oracle view evaluation using the Best-of-N method in Section 2.2.3. For each question, we render images from $k$ viewpoints and sampled $n$ responses per viewpoint. As shown in Figure 5, the result decreases as $n$ increases and converges around $n=20$, indicating that randomness from selecting the maximum score is effectively mitigated. The scores are detailed in Table 4.
Besides the Best-of-N, we explore another method called Human-Intuition-Selection (HIS). We calculate the centroids of all relevant objects of each question and use a heuristic algorithm to select the best viewpoint that captures the centroids effectively. Details and experimental results of these method are in Appendix B.1.
@@
3.4 Human Evaluation
We performed human evaluation on 50 randomly sampled tasks from each benchmark, and use Spearman’s rank correlation (rho) Spearman (1904) to calculate the correlation between original metrics and human judgments, as shown in Table 5.
Benchmark Metric rho
3D MM-Vet LLM evaluation 77.7
ObjaverseXL- Bleu_1 49.8
LVIS Caption ROUGH_L 60.1
CIDEr 50.9
METEOR 28.3
ScanQA ROUGH_L 33.6
CIDEr 25.5
SQA3D EM 80.2
Table 5: The rho between original benchmark metrics and human judgments.
@@ 4 Analysis
4.1 Object Point Cloud
Results demonstrate that VLM significantly outperforms SOTA 3D LLMs using only single view images on these object benchmarks. This suggests that these tasks can be easily solved without specialized 3D representations, indicating a lack of effective 3D capability assessment. The limitation may arise from (1) $\bullet$ the inherent simplicity of these object point clouds, and (2) $\bullet$ these tasks primarily require only surface-level information that could be equally extracted from a simple rendered image. See more task examples in Appendix A.1 and A.2.
@@ 4.2.1 Poor Performance with Single View
In these scene benchmarks, even the most advanced VLMs consistently underperform top 3D models with the single view setting, $\bullet$ revealing their limitations in handling complex 3D scenes.
@@ 4.2.4 Failure of HIS Method
Table 7 shows that HIS performs significantly worse than Best-of-N. Through case studies, we identify three key factors: (1) $\bullet$ Selecting the best viewpoint for detail items in complex scenes with intuition is challenging, and the resulting viewpoints are sometimes suboptimal, primarily due to object occlusions. (2) $\bullet$ Limited ground truth reveals a clear preference. As detailed in Appendix C.2.2, while the viewpoints selected by the HIS method do provide useful information, BoN selects viewpoints that align most closely with the ground truth by definition. As a result, although answers from the HIS method are reasonable, the limited coverage of ground truth makes BoN methods achieve higher scores in some cases. (3) $\bullet$ Many questions rely on common sense, not the specific details of the point cloud, as shown in Appendix C.3. Specific information in HIS viewpoints might even lead to a worse score, compared to answering leveraging world knowledge with an unrelated viewpoint in BoN.
@@ 5 Principles For Effective 3D Evaluation
Based on the analysis, we propose the following principles for effective 3D capabilities evaluation of 3D LLMs.***The analyses leading to each principle are marked with corresponding colored circular markers in the preceding text. We also apply these principles to redesign a task, as shown in Figure 6.
5.1 Principles
@@
5.2 Quantitative Validation
We conduct a quantitative analysis to validate the effectiveness of our principles. The results demonstrate that existing benchmarks indeed fall short of these principles to varying degrees.
LLM Evaluation
We conduct LLM evaluation on the first three principles related to task design. Specifically, we provide GPT-4o with all the queries, ground truth, and rendered point clouds and ask it to determine whether each task complies with these principles.
Pass Rate (%) 3D MM-Vet ObjaverseXL ScanQA SQA3D
-LVIS Caption
Point Cloud Selection 53.9 36.3 69.7 86.2
Task Focus 25.9 19.0 76.5 91.6
Context Specific Inquiry 64.7 93.5 99.2 97.6
All Three Above 16.0 7.3 62.9 81.9
Table 6: The pass rate of each benchmark in LLM Evaluation.
As shown in the Table 6, even the benchmark with the highest pass rate (SQA3D) still has 18.1% of tasks failing to pass all three principles, and the varying pass rates across benchmarks correlate with the performance difference between VLMs and 3D LLMs. For example, lower pass rates are observed on object point cloud benchmarks where VLMs outperform 3D SOTA.
Human Evaluation
Our human evaluation experiments in Section 3.4 strongly support our last principle about evaluation methodology. Using LLM evaluation, 3D MM-Vet shows a high correlation with human judgments . SQA3D achieves an even higher correlation because it’s a multiple-choice task. In contrast, open-ended tasks like ObjaverseXL-LVIS Caption and ScanQA, which rely on text similarity metrics, demonstrate lower correlations. These results support our advocacy for diverse possible answers and flexible LLM-based assessments for open-ended questions.
@@ 7 Limitations
8 Ethics Statement
This research on 3D LLM benchmarks adheres to ethical principles by exclusively utilizing publicly available data and maintaining transparency through open-source code. Our study does not involve any human annotation or privacy concerns. We do not anticipate any potential risks from this methodological study.
@@ Acknowledgement
We would like to express our sincere gratitude to Prof. Yonglu Li for his valuable guidance and support throughout this research, from topic selection to the final writing. His insightful discussions and feedback have been essential to the completion of this work.
@@ C.2.2 ScanQA
Here is an example where the Best-of-N method achieved a significantly better score due to its alignment with ground truth preference: the question is “Where is the brown chair located?” and the ground truths are “under the circular table” and “next to the left table”. The HIS method gives the answer “near the cabinet”, using the input image shown in Figure 14 (a). The answer provided by the Best-of-N method is “on the right side of the table”, based on the input image shown in Figure 14 (b).
It is evident that there are many reasonable answers to this question; however, the ground truth prefers responses related to the table. While both viewpoints depict the chair clearly, the answer from the Human-Intuition-Selection method is influenced by the nearby cabinet in its chosen viewpoint, resulting in a lower score.
(a) The viewpoint selected by Human-Intuition-Selection.
(b) The viewpoint selected by Best-of-N.